# InstructMoLE: Instruction-Guided Mixture of Low-rank Experts for Multi-Conditional Image Generation

### v1

@@ 1 Introduction
The advent of powerful, open-source Diffusion Transformers (DiTs) [peebles2023scalable, esser2024scaling, flux2024] has unlocked unprecedented capabilities in generative AI, fueling a demand for highly specialized and compositional functionalities, from multi-subject composition to personalized content creation [labs2025flux, wu2025qwenimagetechnicalreport, wu2025omnigen2, liu2025step1x-edit]. Parameter-Efficient Fine-Tuning (PEFT), particularly Low-Rank Adaptation (LoRA) [hu2022lora], has become the de facto standard for such customization [xiao2024coap]. However, LoRA’s monolithic update structure conflicts with the demands of multi-task fine-tuning, leading to catastrophic forgetting as different task objectives interfere [biderman2024loralearnsforgets, han2024parameterefficientfinetuninglargemodels].
@@ 3.1.1 Instruction-Guided Routing (IGR) Policy
Here, $\mathcal{I}_{b}\subset\{1,\dots,N\}$ is the set of indices for the Top-$k$ experts, and $\mathbf{w}_{b}\in\mathbb{R}^{k}$ contains their associated, unnormalized $\mathrm{Softmax}$ probabilities. $\mathcal{I}_{b}$ and $\mathbf{w}_{b}$ is then broadcast across all $L$ token positions in the sequence. The final output for the $l$-th layer, $\mathbf{Y}_{b}$, is computed by summing the output of a shared linear layer with the weighted sum of the selected expert outputs:
@@ 4.1 Experimental Setup
We evaluate all models on a suite of benchmarks, each targeting a distinct capability: OmniContext [omnicontext_benchmark_2025] for in-context generation, XVerseBench [chen2025xverse] for multiple subjects-driven generation, GEdit-EN-full [liu2025step1x] for single-image editing, and MultiGen-20M [qin2023unicontrol] and COCO Pose 2017 [lin2014microsoft] for spatially controlled generation.
@@ 4.2 Main Results
On the Complex-Edit benchmark [complexedit_benchmark_2024], which features highly complex, chain-of-thought instructions, InstructMoLE demonstrates robust generalization. It achieves superior perceptual quality (7.15 vs 7.08) and identity preservation (8.09 vs 8.01) compared to Flux.1 Kontext, confirming that the global routing signal maintains high fidelity even when processing complex and potentially ambiguous instructions.
@@ 4.3 Ablation Studies
We conduct a comprehensive evaluation across three key domains. For multiple subjects-driven generation and single-image editing, we report scores on the XVerseBench [chen2025xverse], GEdit-EN-full [gedit_benchmark_2024]. For spatial alignment, we evaluate on 500 samples for each modality, randomly sampled from the evaluation splits of MultiGen-20M [qin2023unicontrol] for depth and Canny edge control, and COCO Pose 2017 [lin2014microsoft] for human pose control. The corresponding metrics are Root Mean Squared Error (RMSE) for depth, F1 score for Canny edges, and an F1 score based on Object Keypoint Similarity (OKS) for pose.
@@ 5 Conclusion
[…]
5.1 The Use of Large Language Models (LLMs)
We acknowledge the use of a large language model (LLM) to aid in the writing process of this manuscript. Its application was confined to language enhancement tasks, such as proofreading for grammatical and spelling errors, rephrasing sentences to improve clarity and readability, and ensuring a consistent formal tone. The LLM served exclusively as a writing assistant.
5.2 Training Data Details
Our model’s versatile capabilities are the direct result of a comprehensive, mixed training dataset, visually summarized in Figure 6. This dataset is meticulously curated to expose the model to a wide spectrum of conditional inputs, combining large-scale, self-synthesized data for complex compositional tasks with established public datasets for foundational spatial control.
Synthesized Data Corpus.
To enable sophisticated, instruction-driven editing and composition, we generated a large-scale corpus of training examples. This corpus focuses on tasks that are poorly represented in public datasets but are crucial for real-world applications. Our synthesis pipeline covers the following key areas:
  • Reference-based Generation: This category involves tasks that condition the output on multiple input images and a textual instruction.
    • Face Swapping: Training triplets consist of a source face image, a target person’s image, and an instruction to transfer the facial identity.
    • Multi-Subject Composition: To teach complex relational reasoning, we generate scenes described by prompts involving multiple, often unrelated, subjects (e.g., “An old man, a French Bulldog, a tuba, and a rattlesnake in a sunny park”).
    • Virtual Try-on: Samples include a person, multiple clothing items, and potentially a pose skeleton, with instructions to dress the person in the specified apparel.
    • Re-Lighting: Data consists of a foreground subject and a new background, with the goal of seamlessly integrating the subject into the new lighting environment.
    • Style Transfer: We provide a content image and a style reference image (e.g., a photograph and a sketch), instructing the model to render the content in the given artistic style.
  • Single Image Editing: We employ the GPT-IMAGE-EDIT-1.5M dataset [wang2025gptimageedit15mmillionscalegptgeneratedimage]. This corpus contains over 1.5 million high-quality samples, constructed by leveraging GPT-4o to unify and refine existing public datasets such as OmniEdit, HQ-Edit, and UltraEdit. The refinement process ensures superior visual fidelity and semantic alignment for instruction-based editing tasks.
Public Datasets.
To build a robust foundation in controllable generation, we incorporate several large-scale, publicly available datasets. These datasets provide strong supervision for fundamental spatial and structural alignment tasks.
  • Spatial Alignment Control: We leverage datasets that provide dense spatial conditioning maps.
    • Depth Control: We utilize datasets such as SubjectSpatial-200K, which provide image-depth map pairs, to train the model to generate scenes with accurate spatial layouts.
    • Canny Edge Control: Similarly, we use datasets with image-Canny edge pairs to enable generation from structural outlines.
    • Pose Control: We use the widely adopted COCO 2017 dataset [lin2014microsoft] with its OpenPose keypoint annotations to teach the model to generate human figures conforming to specific pose skeletons.
  • Image Inpainting (Fill): Using datasets like SubjectSpatial-200K, we generate training examples by randomly masking regions of an image. This trains the model to fill in missing parts coherently, a crucial skill for image editing and out-painting tasks.
Figure 6: Training Data Details.
5.3 Ablation Study
Table 6: Ablation study on the MoLE hyper-parameter configuration.
MoLE Multi-Subject ($\uparrow$) Single-Subject ($\uparrow$) Depth MSE ($\downarrow$) Canny F1 ($\uparrow$) Pose F1 ($\uparrow$)
[…]
rr=32 NN=8 kk=4 65.68 6.15 33.34 41.51% 40.97%
[…]
rr=128 NN=2 kk=1 65.10 6.15 35.50 37.52% 34.02%
Analysis of MoLE Hyper-parameters.
We analyze the MoLE configuration in Table 6 to justify our hyper-parameter choices. The results yield two key insights. First, for a fixed expert pool of $N=8$ and rank $r=32$, performance peaks when activating $k=4$ experts. Activating more experts ($k>4$), especially in the non-sparse case ($k=8$), leads to performance degradation, highlighting the importance of sparse activation in mitigating expert interference. Second, for a fixed activation budget ($r\times k\approx 128$), our configuration with a larger pool of diverse, low-rank experts ($N=8,r=32$) is more effective than alternatives with fewer, high-capacity experts (e.g., $N=4,r=64$). This suggests that a greater diversity of specialized experts is more beneficial than individual expert capacity. These findings validate our use of the ($r=32,N=8,k=4$) configuration for all main experiments.
Analysis of Expert Specialization Dynamics.
Figure 7: Visualization of the expert routing weight distributions in Double Stream Block. The numbers under the "Single Stream Block" column on the left (0, 9, 18) represent the layer index.
Figure 8: Visualization of the expert routing weight distributions in Single Stream Block. The numbers under the "Single Stream Block" column on the left (0, 18, 37) represent the layer index.
To investigate the learning dynamics of our MoE framework, we visualize the expert routing weight distributions at different training stages (1K, 10K, and 100K steps) and across various model depths, as shown in Figure 7 and Figure 8. The analysis reveals two key findings.
First, we observe a clear evolutionary trajectory from an initial, diffuse routing policy to a stable and highly specialized one. At 1K steps, the weights are distributed relatively evenly, indicating an early exploratory phase. By 100K steps, the distributions become notably sparse and peaked, with specific experts consistently chosen for particular task categories. This progression from a generalized to a specialized routing policy demonstrates the effective convergence of our training objectives in guiding functional disentanglement.
Second, the nature of expert specialization varies with model depth, suggesting a hierarchical division of labor. In early layers (e.g., Single Stream Block 0), experts tend to specialize in processing fundamental input types, such as distinguishing “subject" from “objects". In medial layers (e.g., Single Stream Block 18), we observe the emergence of clear, task-level specialists, with distinct experts strongly favoring tasks like “canny" or “swapface". In later layers (e.g., Single Stream Block 37), the routing pattern often becomes more distributed again, suggesting a shift from semantic task execution to a more collaborative final synthesis stage involving multiple experts.
5.4 Qualitative Comparison
Figure 9: Qualitative comparison on OmniContext benchmark.
Figure 10: Qualitative comparison on Xverse benchmark.
Figure 11: Sample outputs generated by our model (InstructMoLE) from diverse, multi-modal instructions.
5.5 Qualitative Analysis of Extreme Scenarios
Figure 12: Extreme Scenario Stress Test: Flux.1 Kontext vs. InstructMoLE. We evaluate robustness across five edge cases spanning input noise, implicit reasoning, and complex composition. While both models share limitations under severe textual corruption (#1) and high reference load (#5), InstructMoLE demonstrates superior capabilities in abstract atmospheric inference (#2) and logical causality (#4), successfully executing implicit edits where the baseline fails.
To rigorously assess the operational boundaries of our method, we conducted stress tests on extreme scenarios involving textual noise, implicit reasoning, and high-complexity composition, as visualized in Figure 12.
Limitations and Common Failure Modes. Our analysis reveals a discernible threshold for robustness. In Case #1, characterized by severe typos (e.g., “Chnage the bckgrnd…”), both InstructMoLE and the baseline fail to resolve the semantic intent. This indicates that while the CLIP semantic anchor provides stability, it cannot fully compensate for T5 token embeddings when textual corruption is excessive. Similarly, in the ultra-complex Case #5, which requires simultaneous reference binding for three distinct entities (“robot,” “boy,” “girl”), both models struggle to correctly map identities to the generated subjects. This points to a capacity bottleneck in the underlying cross-attention mechanism of the foundation model when handling high reference loads, a limitation that persists regardless of the routing strategy.
Superiority in Reasoning and Coherence. Despite these boundary conditions, InstructMoLE exhibits significantly stronger capabilities in implicit reasoning and atmospheric consistency. In Case #2 (“Make it look dangerous”), our model successfully modulates the robot’s expression and lowers scene lighting to align with the semantic tone, whereas the baseline output remains incongruously bright. Crucially, in Case #4 (“The person is cold”), InstructMoLE correctly infers the causal implication to generate clothing (a jacket), a logical step the baseline fails to execute. Furthermore, in multi-attribute editing (Case #3), our method achieves superior spatial and lighting integration of the inserted object.

### v2

@@ 1 Introduction
The advent of powerful, open-source Diffusion Transformers (DiTs) [peebles2023scalable, esser2024scaling, flux2024] has unlocked unprecedented capabilities in generative AI, fueling a demand for highly specialized and compositional functionalities, from multi-subject composition to personalized content creation [labs2025flux, wu2025qwenimagetechnicalreport, wu2025omnigen2, liu2025step1xeditpracticalframeworkgeneral]. Parameter-Efficient Fine-Tuning (PEFT), particularly Low-Rank Adaptation (LoRA) [hu2022lora], has become the de facto standard for such customization [xiao2024coap]. However, LoRA’s monolithic update structure conflicts with the demands of multi-task fine-tuning, leading to catastrophic forgetting as different task objectives interfere [biderman2024loralearnsforgets, han2024parameterefficientfinetuninglargemodels].
@@ 3.1.1 Instruction-Guided Routing (IGR) Policy
Here, $\mathcal{I}_{b}\subset\{1,\dots,N\}$ is the set of indices for the Top-$k$ experts, and $\mathbf{w}_{b}\in\mathbb{R}^{k}$ contains their associated, unnormalized $\mathrm{Softmax}$ probabilities. $\mathcal{I}_{b}$ and $\mathbf{w}_{b}$ are then broadcast across all $L$ token positions in the sequence. The final output for the $l$-th layer, $\mathbf{Y}_{b}$, is computed by summing the output of a shared linear layer with the weighted sum of the selected expert outputs:
@@ 4.1 Experimental Setup
We evaluate all models on a suite of benchmarks, each targeting a distinct capability: OmniContext [wu2025omnigen2] for in-context generation, XVerseBench [chen2025xverse] for multiple subjects-driven generation, GEdit-EN-full [liu2025step1xeditpracticalframeworkgeneral] for single-image editing, and MultiGen-20M [qin2023unicontrol] and COCO Pose 2017 [lin2014microsoft] for spatially controlled generation.
@@ 4.2 Main Results
On the Complex-Edit benchmark [yang2025textttcomplexeditcotlikeinstructiongeneration], which features highly complex, chain-of-thought instructions, InstructMoLE demonstrates robust generalization. It achieves superior perceptual quality (7.15 vs 7.08) and identity preservation (8.09 vs 8.01) compared to Flux.1 Kontext, confirming that the global routing signal maintains high fidelity even when processing complex and potentially ambiguous instructions.
@@ 4.3 Ablation Studies
We conduct a comprehensive evaluation across three key domains. For multiple subjects-driven generation and single-image editing, we report scores on the XVerseBench [chen2025xverse], GEdit-EN-full [liu2025step1xeditpracticalframeworkgeneral]. For spatial alignment, we evaluate on 500 samples for each modality, randomly sampled from the evaluation splits of MultiGen-20M [qin2023unicontrol] for depth and Canny edge control, and COCO Pose 2017 [lin2014microsoft] for human pose control. The corresponding metrics are Root Mean Squared Error (RMSE) for depth, F1 score for Canny edges, and an F1 score based on Object Keypoint Similarity (OKS) for pose.